OpenStudy (anonymous):

help on solving x^2n-2x^n-3 Most of the time people say you can't solve it but my math book says the answer is (x^n-3)(x^n+1). I understand how to factor it but... what happened to the x^2n? The 2? Because once you put it in this form after you x-factor: (x^2n-3x^n+1x^n-3) and you group them (x^2n-3x^n)+(1x^n-3) When you take the common factor out to simplify you get: X^n(x^2-3)+1(x^n-3) ....am i doing something wrong?

OpenStudy (amistre64):

x^(2n) = (x^n)^2

OpenStudy (amistre64):

id solve for u^2 -2u -3, given that u=x^n

OpenStudy (mathmale):

x^2n-2x^n-3 is ambiguous, and should be written with parentheses as amistre64 has show you here. It never hurts to use parentheses, whereas leaving out necessary parentheses can lead to great confusion.

OpenStudy (mathmale):

u^2 -2u -3 is a quadratic expression. Which methods of factoring quadratics do you know? Apply one here.

OpenStudy (anonymous):

i know the -b+-Radical(b)^2-4ac / 2a

OpenStudy (amistre64):

x^n (x^2-3) x^n x^2 is not equal to x^(2n) x^n x^2 =x^(2+n)

OpenStudy (amistre64):

x^n x^n = x^(n+n) = x^(2n)

OpenStudy (anonymous):

oh i see

OpenStudy (anonymous):

so how do i take it out of that final step after you group them?

OpenStudy (anonymous):

i just take one x^n out?

OpenStudy (amistre64):

factor it correctly :)

OpenStudy (anonymous):

oh oh!!!!! so if i take only one x^2n out i get X^n left still!

OpenStudy (anonymous):

SHHHHHH THANK YOU

OpenStudy (amistre64):

yep

OpenStudy (amistre64):

(x^n x^n-3x^n)+(x^n-3) x^n (x^n-3)+(x^n-3) etc...